Meeting Notes for 7 December 2011 
Present: Phong, TomL, Bob, Liz, TomJ
Absent: TomM, Marge, KenC
  Action Items 
 
-  Set up a TC for CMS calibration (TomM/Phong/Mihael?/TomL)
 
 
  Cosmic 
 
-  Whee blessing in progress
  -  Shower fix being tested in production
  -  APS meeting on 100 years of Cosmic Rays in April 2012
  -  Coordinated data taking - tie into 100 years? 
 
 
  CMS 
 
-  Muon TT/GG/TG data being pushed into the eLab - hope to have by this Friday in the system as a prototype 
  -  Annoying MySQL bugs cropped up, still need to move stuff over to Postgres
  -  Calibration?  
-  Rip off the band-aid and do things right. 
 
 
 
 
  LIGO 
 
-  Mihael will try and fix some bugs during his break (including some data import issues) 
 
 
-- Main.PhongNguyen - 07 Dec 2011